OpenStudy (anonymous):

I need help with this. Explain (in your own words) the difference between zero divided by seven and seven divided by zero. Show a way of writing these two problems using mathematical symbols for division. Don’t forget to write the solution to each one! Do you have an easy way to remember the difference between the two?

Directrix (directrix):

Do you have an easy way to remember the difference between the two? ----- Not entirely good mathematics, but here goes: When zero is underfound, it is undefined. This means that if zero is found in the denominator the result is undefined as in 7/0 is undefined.

OpenStudy (anonymous):

\[(n)\div( 0 ) =\infty \]

OpenStudy (anonymous):

(n)÷(0)=∞ where n =0,1,2,3,..... \[\infty\]

OpenStudy (anonymous):

And \[0\div (n)\]\[=0\] where n n =0,1,2,3,..... ∞